OpenStudy (anonymous):

Solve: 7(x - 1) = 5(x + 2) + 1

OpenStudy (anonymous):

7x-7 = 5x + 10 +1 can you do it now ?

OpenStudy (anonymous):

7x-7=5x+10+1 2x-7=11 2x=18 x=8

OpenStudy (anonymous):

This is wrong

OpenStudy (anonymous):

oh its 9

OpenStudy (anonymous):

it's 9 you should let people do something by thier own

OpenStudy (anonymous):

their*